# Constrained Stochastic Recursive Momentum Successive Convex Approximation

### v3

@@ Abstract
We consider stochastic optimization problems with non-convex functional constraints, such as those arising in trajectory generation, sparse approximation, and robust classification. To this end, we put forth a recursive momentum-based accelerated successive convex approximation (SCA) algorithm. At each iteration, the proposed algorithm entails constructing convex surrogates of the stochastic objective and the constraint functions, and solving the resulting convex optimization problem. A recursive update rule is employed to track the gradient of the stochastic objective function, which contributes to variance reduction and hence accelerates the algorithm convergence. A key ingredient of the proof is a new parameterized version of the standard Mangasarian-Fromowitz Constraints Qualification, that allows us to bound the dual variables and hence obtain problem-dependent bounds on the rate at which the iterates approach an $\epsilon$-stationary point. Remarkably, the proposed algorithm achieves near-optimal stochastic first-order (SFO) complexity with adaptive step sizes and optimal SFO complexity with non-adaptive step sizes at par with that achieved by state-of-the-art stochastic optimization algorithms for solving unconstrained problems. As an example, we detail a obstacle-avoiding trajectory optimization problem that can be solved using the proposed algorithm and show that its performance is superior to that of the existing algorithms used for trajectory optimization. The performance of the proposed algorithm is also shown to be comparable to that of a specialized sparse classification algorithm applied to a binary classification problem.
@@ I-A Background
Existing approaches to solve ($\mathcal{P}$) include the primal-dual method [10], penalty method [11], proximal point methods [12, 13], augmented Lagrangian method [14], and several Successive Convex Approximation (SCA) variants [15, 16, 17, 18, 19]. The performance of these algorithms is usually measured in terms of their oracle complexity, which is the number of calls to the stochastic first-order (SFO) oracle required to reach a point that satisfies the Karush-Kuhn Tucker (KKT) conditions to within a tolerance of $\epsilon$. The primal-dual method proposed in [10] iteratively updates both primal and dual variables based on a stochastic approximation to augmented Lagrangian function, achieving an oracle complexity of $\mathcal{O}(\epsilon^{-2.5})$. A rate of $\mathcal{O}(\epsilon^{-3.5})$ is achieved by [11] which utilizes a penalty functional to reformulate ($\mathcal{P}$) as an unconstrained optimization problem. An improved rate of $\mathcal{O}(\epsilon^{-2})$ is achieved by the Inexact Constrained Proximal Point Algorithm (ICPPA) [13] and level-constrained stochastic proximal gradient (LCSPG) methods [12], both of which solve a sequence of strongly convex subproblems, obtained by adding a quadratic term to the objective and constraint functions. Both [13] and [12] achieve $\mathcal{O}(\epsilon^{-2})$ SFO complexity using fixed $\epsilon$-dependent step-sizes under a standard Lipschitz smoothness assumption on $U$. A related level-constrained proximal point (LCPP) method [9] also achieved the a rate of ${\mathcal{O}}(\epsilon^{-2})$ but was customized to only handle sparsity inducing constraint functions. Improving on these works, the augmented Lagrangian method (MLALM) [14] requires a weaker mean-squared smoothness (MSS) assumption but still achieves the optimal rate of $\mathcal{O}(\epsilon^{-3/2})$, albeit for a fixed $\epsilon$-dependent step-size. In contrast, modern unconstrained stochastic optimization methods such as STORM and its variants have moved to using adaptive step-sizes that are generally easier to tune [20].
@@
We remark that while the oracle complexity of ${\mathcal{O}}(\epsilon^{-2})$ achieved by various algorithms that solve ($\mathcal{P}$) may be optimum under the Lipschitz smoothness assumption on $U$, it can be improved to $\mathcal{O}(\epsilon^{-3/2})$ under the MSS assumption on $f$ [29, (4)]. Indeed, using momentum or variance reduction techniques [30, 31, 32, 33, 34], it is well-known that the optimal oracle complexity of $\mathcal{O}(\epsilon^{-3/2})$ can be achieved for unconstrained or convex-constrained version of ($\mathcal{P}$). Interestingly, variance reduction is applied to the level-constrained proximal algorithm in [12] for finite-sum problems but the rate achieved is only ${\mathcal{O}}(\epsilon^{-2})$, with a remark that it can be improved to $\mathcal{O}(\epsilon^{-3/2})$ if $f({\mathbf{x}},{\boldsymbol{\xi}})$ is Lipschitz smooth in ${\mathbf{x}}$ for all ${\boldsymbol{\xi}}$, but without any proof [12, Remark 5]. As noted earlier, only the recently published work [14] achieves the optimal rate of $\mathcal{O}(\epsilon^{-3/2})$, although under a restrictive condition requiring a fixed $\epsilon$-dependent step size.
@@ I-B Contributions
  • Accelerated Stochastic SCA algorithm: we incorporate the STOchastic Recursive Momentum (STORM) updates from [20, 30] into the SCA framework to achieve near-optimal SFO complexity under adaptive step sizes and optimal SFO complexity under non-adaptive step sizes. Drawing inspiration from these works, the proposed Constrained STORM SCA (CoSTA) algorithm leverages momentum-enhanced updates to improve convergence rates.
@@
${\tilde{\mathcal{O}}}(\epsilon^{-3/2})$
${\mathcal{O}}(\epsilon^{-3/2})$
Parameterized MFCQ
@@
SCA-based, ${\tilde{\mathcal{O}}}(\epsilon^{-3/2})$ with adaptive step-size
and ${\mathcal{O}}(\epsilon^{-3/2})$ with non-adaptive step-size
@@ II Motivating Examples
In this work, we consider the problem of designing the energy-optimal trajectories of $N$ agents, while avoiding an obstacle. Mathematically, the problem entails finding a collection of $T$ waypoints $\{{\mathbf{x}}_{i}(\tau)\}_{\tau=1}^{T}$ for each agent $i$, collected in the super-vector $\underline{{\mathbf{x}}}\in{\mathbb{R}}^{2NT}$. The waypoints are equally spaced at intervals of $\Delta t=T_{f}/T$ seconds where $T_{f}$ is the given time horizon. The $i$-th agent starts at a given location ${\mathbf{x}}_{i}(0)$ and wants to reach the goal location ${\mathbf{x}}_{i}^{g}$, such that ${\mathbf{x}}_{i}(T)={\mathbf{x}}_{i}^{g}$. For collision avoidance, we assume that the agents and the obstacle have a circular shape with radii $r$ and $r^{o}$, respectively. Finally, consider the agent’s movement from ${\mathbf{x}}_{i}(\tau)$ to ${\mathbf{x}}_{i}(\tau+1)$, where $10\leq\tau\leq T-$ 1. The control input — and consequently, the energy — can be considered proportional to the quantity $\left\|{\mathbf{x}}_{i}(\tau+1)-{\mathbf{x}}_{i}(\tau)-\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\Delta t\right\|^{2}$. Here, the term $\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\Delta t$ approximates the vehicle’s displacement without any control input. Hence, the vehicle should seek to minimize the cumulative energy input while ensuring that the magnitude of the control input does not exceed the vehicle specifications, denoted by $v_{i}^{\max}$. In summary, the considered problem can be written as
@@
$\left\|{\mathbf{x}}_{i}(\tau+1)-{\mathbf{x}}_{i}(\tau)-\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\Delta t\right\|\leq v_{i}^{\max}\Delta t\hskip 5.69054pt\forall~{}\tau,i,{\boldsymbol{\xi}}$ (1e)
where the objective seeks to minimize the cumulative energy consumption averaged over the given ensembles, (1b) enforces the termination condition, (1c)-(1d) ensure that the vehicles maintain safe distances from each other and from the obstacle, and (1e) limits the maximum control input for every agent, time, and realization. Of these, (1e) contains an impractically large number of constraints, one for each realization, and renders the problem difficult to solve. To avoid this issue, we consider a more stringent reformulation of the constraint that implies (1e). To this end, we assume that the ensemble forecasts have a uniformly bounded support, i.e., $\left\|\vartheta({\mathbf{x}},{\boldsymbol{\xi}})-{\mathbb{E}}\left[\vartheta({\mathbf{x}},{\boldsymbol{\xi}})\right]\right\|\leq\Delta\vartheta^{\max}$ for ${\mathbf{x}}\in{\mathcal{B}}$. Introducing ${\mathbb{E}}\left[\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\right]\Delta t$ inside the norm on the left of (1e) and using the triangle inequality, we observe that
@@
$\left\|{\mathbf{x}}_{i}(\tau+1)-{\mathbf{x}}_{i}(\tau)-{\mathbb{E}}\left[\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\right]\Delta t\right\|\leq(v_{i}^{\max}-\Delta\vartheta^{\max})\Delta t$ (3)
@@ III-A Problem
$\min_{{\mathbf{x}}\in{\mathcal{K}}}$ $~{}~{}{\mathbb{E}}[f({\mathbf{x}},{\boldsymbol{\xi}})]+u({\mathbf{x}})$ s. t. $~{}~{}~{}~{}{\mathbf{g}}({\mathbf{x}})\leq 0$ ($\mathcal{P^{\prime}}$)
@@ III-B Proposed Algorithm
We now describe the proposed successive convex optimization algorithm in detail. The algorithm is initialized at an arbitrary feasible point ${\mathbf{x}}_{1}={\mathbf{x}}_{0}\in{\mathcal{K}}$ that satisfies ${\mathbf{g}}({\mathbf{x}}_{1})\leq 0$. At iteration $t$ and given iterate ${\mathbf{x}}_{t}$, we construct a strongly convex and smooth surrogate ${\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ of the objective function $f({\mathbf{x}},{\boldsymbol{\xi}}_{t})$ as well as smooth and convex surrogates $\{{\tilde{g}}_{j}({\mathbf{x}},{\mathbf{x}}_{t})\}_{j=1}^{J}$ (collected into the vector-valued function ${\tilde{{\mathbf{g}}}}({\mathbf{x}},{\mathbf{x}}_{t})$) of the constraint functions $\{g_{j}({\mathbf{x}})\}_{j=1}^{J}$. These surrogates are subsequently used to formulate the following convex sub-problem that must be solved at every iteration:
@@
${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t}):={\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+(1-\beta_{t})\langle{\mathbf{x}}-{\mathbf{x}}_{t},{\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t})\rangle$ (6)
@@
We can view ${\mathbf{z}}_{t+1}$ as a convex combination of the unbiased gradient $\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ and the variance-reduced gradient estimate ${\mathbf{z}}_{t}+\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t})$ of SARAH [44], similar to the STORM update in [20] but different from the classical tracking rules used in [27, 15, 45]. Substituting (7) in (6), we see that ${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})={\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\langle{\mathbf{x}}-{\mathbf{x}}_{t},{\mathbf{z}}_{t+1}-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})\rangle$. That is, the running approximation ${\tilde{f}}$ comprises of the surrogate ${\hat{f}}$ and an additional linear correction term that depends on the gradient tracking error. If the surrogate follows the tangent matching property, i.e., $\nabla{\hat{f}}({\mathbf{x}},{\mathbf{x}},{\boldsymbol{\xi}})=\nabla f({\mathbf{x}},{\boldsymbol{\xi}})$, then the update in (7) also implies that
$\nabla{\tilde{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})$ $=\nabla{\hat{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+(1-\beta_{t})({\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t}))$ (8)
@@
for $\eta_{t}<1$.
The choice of the step-size and momentum rules are similar to those in [20]. Specifically, define $G_{t}:=\left\|\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})\right\|$ so that

### v4

@@ Abstract
We consider stochastic optimization problems with non-convex functional constraints, such as those arising in trajectory generation, sparse approximation, and robust classification. To this end, we put forth a recursive momentum-based accelerated successive convex approximation (SCA) algorithm. At each iteration, the proposed algorithm entails constructing convex surrogates of the stochastic objective and the constraint functions, and solving the resulting convex optimization problem. A recursive update rule is employed to track the gradient of the stochastic objective function, which contributes to variance reduction and hence accelerates the algorithm convergence. A key ingredient of the proof is a new parameterized version of the standard Mangasarian-Fromowitz Constraints Qualification, that allows us to bound the dual variables and hence obtain problem-dependent bounds on the rate at which the iterates approach an $\epsilon$-stationary point. Remarkably, the proposed algorithm achieves near-optimal stochastic first-order (SFO) complexity with adaptive step sizes closely matching that achieved by state-of-the-art stochastic optimization algorithms for solving unconstrained problems. As an example, we detail an obstacle-avoiding trajectory optimization problem that can be solved using the proposed algorithm and show that its performance is superior to that of the existing algorithms used for trajectory optimization. The performance of the proposed algorithm is also shown to be comparable to that of a specialized sparse classification algorithm applied to a binary classification problem.
@@ I-A Background
Existing approaches to solve ($\mathcal{P}$) include the primal-dual method [10], penalty method [11], proximal point methods [12, 13], augmented Lagrangian method [14], and several Successive Convex Approximation (SCA) variants [15, 16, 17, 18, 19]. The performance of these algorithms is usually measured in terms of their oracle complexity, which is the number of calls to the stochastic first-order (SFO) oracle required to reach a point that satisfies the Karush-Kuhn Tucker (KKT) conditions to within a tolerance of $\epsilon$. The primal-dual method proposed in [10] iteratively updates both primal and dual variables based on a stochastic approximation to augmented Lagrangian function, achieving an oracle complexity of $\mathcal{O}(\epsilon^{-2.5})$. A rate of $\mathcal{O}(\epsilon^{-3.5})$ is achieved by [11] which utilizes a penalty functional to reformulate ($\mathcal{P}$) as an unconstrained optimization problem. An improved rate of $\mathcal{O}(\epsilon^{-2})$ is achieved by the Inexact Constrained Proximal Point Algorithm (ICPPA) [13] and level-constrained stochastic proximal gradient (LCSPG) methods [12], both of which solve a sequence of strongly convex subproblems, obtained by adding a quadratic term to the objective and constraint functions. Both [13] and [12] achieve $\mathcal{O}(\epsilon^{-2})$ SFO complexity using $\epsilon$-dependent step-sizes under a standard Lipschitz smoothness assumption on $U$. A related level-constrained proximal point (LCPP) method [9] also achieved the a rate of ${\mathcal{O}}(\epsilon^{-2})$ but was customized to only handle sparsity inducing constraint functions. Improving on these works, the augmented Lagrangian method (MLALM) [14] requires a stronger mean-squared smoothness (MSS) assumption and achieves the optimal rate of $\mathcal{O}(\epsilon^{-3/2})$, though for a fixed $\epsilon$-dependent step-size. In contrast, modern unconstrained stochastic optimization methods such as STORM and its variants have moved to using adaptive step-sizes that are generally easier to tune [20], albeit under assumptions even stronger than MSS.
@@
We remark that while the oracle complexity of ${\mathcal{O}}(\epsilon^{-2})$ achieved by various algorithms that solve ($\mathcal{P}$) may be optimum under the Lipschitz smoothness assumption on $U$, it can be improved to $\mathcal{O}(\epsilon^{-3/2})$ under the MSS assumption on $f$ [29, (4)]. Indeed, using momentum or variance reduction techniques [30, 31, 32, 33, 34], it is well-known that the optimal oracle complexity of $\mathcal{O}(\epsilon^{-3/2})$ can be achieved for unconstrained or convex-constrained version of ($\mathcal{P}$). Interestingly, variance reduction is applied to the level-constrained proximal algorithm in [12] for finite-sum problems but the rate achieved is only ${\mathcal{O}}(\epsilon^{-2})$, with a remark that it can be improved to $\mathcal{O}(\epsilon^{-3/2})$ if $f({\mathbf{x}},{\boldsymbol{\xi}})$ is Lipschitz smooth in ${\mathbf{x}}$ for all ${\boldsymbol{\xi}}$, but without any proof [12, Remark 5]. As noted earlier, only the recently published work [14] achieves the optimal rate of $\mathcal{O}(\epsilon^{-3/2})$ with MSS assumption, although under a restrictive condition requiring a fixed $\epsilon$-dependent step size.
@@ I-B Contributions
  • Accelerated Stochastic SCA algorithm: we incorporate the STOchastic Recursive Momentum (STORM) updates from [20, 30] into the SCA framework to achieve near-optimal SFO complexity with non-adaptive step sizes under the MSS assumption and with adaptive step sizes under a stronger smoothness assumption. Drawing inspiration from these works, the proposed Constrained STORM SCA (CoSTA) algorithm leverages momentum-enhanced updates to improve convergence rates.
@@
${\tilde{\mathcal{O}}}(\epsilon^{-3/2})$ Parameterized MFCQ
@@
SCA-based, non-adaptive step-size
requires only MSS assumption
@@ II Motivating Examples
We consider the problem of designing the energy-optimal trajectories of $N$ agents, while avoiding an obstacle. Mathematically, the problem entails finding a collection of $T$ waypoints $\{{\mathbf{x}}_{i}(\tau)\}_{\tau=1}^{T}$ for each agent $i$, collected in the super-vector $\underline{{\mathbf{x}}}\in{\mathbb{R}}^{2NT}$. The waypoints are equally spaced at intervals of $\Delta t=T_{f}/T$ seconds where $T_{f}$ is the given time horizon. The $i$-th agent starts at a given location ${\mathbf{x}}_{i}(0)$ and wants to reach the goal location ${\mathbf{x}}_{i}^{g}$, such that ${\mathbf{x}}_{i}(T)={\mathbf{x}}_{i}^{g}$. For collision avoidance, we assume that the agents and the obstacle have a circular shape with radii $r$ and $r^{o}$, respectively. Finally, consider the agent’s movement from ${\mathbf{x}}_{i}(\tau)$ to ${\mathbf{x}}_{i}(\tau+1)$, where $0\leq\tau\leq T-1$. The control input — and consequently, the energy — can be considered proportional to the quantity $\left\|{\mathbf{x}}_{i}(\tau+1)-{\mathbf{x}}_{i}(\tau)-\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\Delta t\right\|^{2}$. Here, the term $\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\Delta t$ approximates the vehicle’s displacement without any control input. Hence, the vehicle should seek to minimize the cumulative energy input while ensuring that the magnitude of the control input does not exceed the vehicle specifications, denoted by $v_{i}^{\max}$. In summary, the considered problem can be written as
@@
$\hskip-19.91692pt\left\|{\mathbf{x}}_{i}(\tau+1)-{\mathbf{x}}_{i}(\tau)-\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\Delta t\right\|\leq v_{i}^{\max}\Delta t\hskip 5.69054pt\forall~\tau,i,{\boldsymbol{\xi}}$ (1e)
where the objective seeks to minimize the cumulative energy consumption averaged over the given ensembles, (1b) enforces the termination condition, (1c)-(1d) ensure that the vehicles maintain safe distances from each other and from the obstacle, and (1e) limits the maximum control input for every agent, time, and realization. Of these, (1e) contains an impractically large number of constraints, one for each realization, and renders the problem difficult to solve. To avoid this issue, we consider a more stringent reformulation of the constraint, i.e., (1e). To this end, we assume that the ensemble forecasts have a uniformly bounded support, i.e., $\left\|\vartheta({\mathbf{x}},{\boldsymbol{\xi}})-{\mathbb{E}}\left[\vartheta({\mathbf{x}},{\boldsymbol{\xi}})\right]\right\|\leq\Delta\vartheta^{\max}$ for ${\mathbf{x}}\in{\mathcal{B}}$. Introducing ${\mathbb{E}}\left[\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\right]\Delta t$ inside the norm on the left of (1e) and using the triangle inequality, we observe that
@@
$\left\|{\mathbf{x}}_{i}(\tau+1)-{\mathbf{x}}_{i}(\tau)-{\mathbb{E}}\left[\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\right]\Delta t\right\|$
$\hskip-39.83368pt\leq(v_{i}^{\max}-\Delta\vartheta^{\max})\Delta t$ (3)
@@ III-A Problem
$\min_{{\mathbf{x}}\in{\mathcal{K}}}$ $~~{\mathbb{E}}[f({\mathbf{x}},{\boldsymbol{\xi}})]+u({\mathbf{x}})$ s. t. $~~~~{\mathbf{g}}({\mathbf{x}})\leq 0.$ ($\mathcal{P^{\prime}}$)
@@ III-B Proposed Algorithm
We now describe the proposed SCA algorithm in detail. The algorithm is initialized at an arbitrary feasible point ${\mathbf{x}}_{1}={\mathbf{x}}_{0}\in{\mathcal{K}}$ that satisfies ${\mathbf{g}}({\mathbf{x}}_{1})\leq 0$. At iteration $t$ and given iterate ${\mathbf{x}}_{t}$, we construct a strongly convex and smooth surrogate ${\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ of the objective function $f({\mathbf{x}},{\boldsymbol{\xi}}_{t})$ as well as smooth and convex surrogates $\{{\tilde{g}}_{j}({\mathbf{x}},{\mathbf{x}}_{t})\}_{j=1}^{J}$ (collected into the vector-valued function ${\tilde{{\mathbf{g}}}}({\mathbf{x}},{\mathbf{x}}_{t})$) of the constraint functions $\{g_{j}({\mathbf{x}})\}_{j=1}^{J}$. These surrogates are subsequently used to formulate the following convex sub-problem that must be solved at every iteration:
@@
$\hskip-85.35826pt{\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t}):={\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ (6)
$\hskip-71.13188pt+(1-\beta_{t})\langle{\mathbf{x}}-{\mathbf{x}}_{t},{\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t})\rangle$
@@
We can view ${\mathbf{z}}_{t+1}$ as a convex combination of the unbiased gradient $\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ and the variance-reduced gradient estimate ${\mathbf{z}}_{t}+\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t})$ of SARAH [44], similar to the STORM update in [20] but different from the classical tracking rules used in [27, 15, 45]. Substituting (7) in (6), we see that ${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})={\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\langle{\mathbf{x}}-{\mathbf{x}}_{t},{\mathbf{z}}_{t+1}-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})\rangle$. That is, the running approximation ${\tilde{f}}$ comprises of the surrogate $^{\hat{f}}$ and an additional linear correction term that depends on the gradient tracking error. If the surrogate follows the tangent matching property, i.e., $\nabla{\hat{f}}({\mathbf{x}},{\mathbf{x}},{\boldsymbol{\xi}})=\nabla f({\mathbf{x}},{\boldsymbol{\xi}})$, then the update in (7) also implies that
$\nabla{\tilde{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})$
$=\nabla{\hat{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+(1-\beta_{t})({\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t}))$ (8)
@@
for $\eta_{t}<1$. The choice of the step-size and momentum rules are similar to those in [20]. Specifically, define $G_{t}:=\left\|\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})\right\|$ so that